OpenStudy (anonymous):

The compound interest on a certain sum equals to Rs 510 in 2 years at 4 per cent annual rate of interest. Find the simple interest on the same amount for the same period if the rate of interest is double that given for. compound interest?

OpenStudy (anonymous):

Simple interest=P*n*r Make data that will help you alot :) Data: P=Principle=510 n=number of years=2 r=rate=4%(BUT NOW ITS SAYING THAT IT GOT DOUBLED SO THE RATE OF INTEREST WILL BE)=8%(change it in decimal:8/100=0.08 put the values in the equation Simple interest=P*n*r Simple interest=510*2*0.08 Simple interest=81.6 Do you any options so just lemme know I'll see if I can help you again :)